# Rethinking MUSHRA: Addressing Modern Challenges in Text-to-Speech Evaluation

### v1

@@ 6.2 Evaluations using MUSHRA-DG
Does our proposed variant help mitigate judgement ambiguity while rating? In Table 4, we show the effects of presenting detailed guidelines along with scoresheets to 14 participants to systematically arrive at MUSHRA scores. We find MUSHRA-DG scores to be rank-consistent with the CMOS tests , while systems scores are much higher and closer to the “Excellent" label, as expected. More importantly, the standard deviation of scores across all systems reduced by 41% in Hindi and 58% in Tamil when compared to the original MUSHRA, indicating that our proposed variant is able to reduce the ambiguity of rating naturalness on a single bar while preserving ranks.
@@
Time Complexity of MUSHRA-DG. We hypothesize that the additional detail of evaluating each audio sample across multiple dimensions inevitably increases the time required for participants to complete the test. To verify this hypothesis, we visualize the average time taken across pages in Figure 6 and find that the MUSHRA-DG test indeed takes nearly twice as much time as the original MUSHRA test. However, we believe this extra time results in a much more comprehensive understanding of TTS system performance, making the trade-off worthwhile.
@@ 7 Conclusion
Our comprehensive study reveals significant shortcomings in the current use of the MUSHRA test for evaluating modern high-quality TTS systems. Through an extensive analysis involving 47,100 human ratings, we identified two primary issues: reference-matching bias and judgment ambiguity. To address these issues, we propose two refined variants of the MUSHRA test: MUSHRA-NMR, which omits explicit identification of the human reference, and MUSHRA-DG, which uses detailed guidelines to calculate MUSHRA scores systematically. Our findings indicate that both variants lead to more reliable evaluations, with MUSHRA-DG offering the additional benefit of fine-grained fault isolation during assessment. Through this work, we also release MANGO, a large human rating dataset, to further support research in this area.
@@ A.2 Sensitivity of MUSHRA in Tamil
Figure 8: Spearman rank correlation of MUSHRA scores in Tamil using subsets of listeners and utterances vs. scores of all listeners and utterances.
A.3 Detailed Guidelines for MUSHRA-DG
@@
$s_{M}$ $=\frac{L+VQ+R}{3}$
$\quad-\min(MP,15)\times 5$
$\quad-\min(7,SP)\times 10$
$\quad-US\times 5$
$\quad-DA\times 5$
$\quad-WS\times 25$
$\quad-SEF\times 5$
A.4 MUSHRA-Extended
Table 5: MUSHRA-Extended scores with 95% CI for Hindi.
System $\mu$ $\sigma$
CI
FS2 63.12 21.3 0.93
ST2 65.15 21.76 0.95
VITS-R 68.47 19.70 0.86
VITS 68.99 19.67 0.86
ANC 73.62 19.56 0.79
XTTS 73.65 18.52 0.86
REF 76.39 18.05 0.81
A.5 Visualizing MUSHRA Distributions
@@
Figure 12: Visualization of the MUSHRA score distributions per utterance across three systems— FS2, ST2, and VITS, along with the reference (REF) and anchor (ANC) for Tamil. The X-axis represents each of the 100 utterances. Each boxplot represents ratings (0-100) across all raters for a system for a given utterance. The substantial heights of some boxplots indicate significant variance in the scores given by different raters for a single utterance.
A.6 Limitations
Our study focuses on human evaluations for Hindi and Tamil, representing a major Indo-Aryan and Dravidian language, respectively. However, we did not extend our analysis to English, a widely spoken and diverse language. This limitation is due to the scope of our current research and resource constraints. Future studies should include evaluations in English to generalize our findings across different language families and understand how language-specific characteristics might influence TTS evaluation outcomes. This broader analysis could provide more comprehensive insights into the applicability and robustness of our proposed MUSHRA variants across diverse linguistic contexts.
A.7 Ethical Considerations
We prioritized ethical conduct throughout our research. The 471 human listeners involved in the study provided informed consent before participating in the evaluation, recruited through professional data annotation agencies. These agencies verified participant language proficiency for task relevance. We established an education criterion of completing grade 12 (Indian system) to ensure participants’ ability to accurately annotate audio content. Participants were compensated fairly for their time and expertise, following industry standards. They were also fully informed about the study nature, procedures, and their right to withdraw at any point without consequence.
We release the evaluations dataset, which includes 47,100 human ratings, under CC-BY-4.0 license after careful consideration of privacy and ethical use. Identifiable information about the participants was anonymized to protect their privacy. We encourage the use of this dataset for advancing TTS evaluation metrics, emphasizing that it should be used responsibly and ethically, adhering to principles of transparency and fairness. Finally, we acknowledge that our study focuses on Hindi and Tamil, and we recognize the importance of extending such evaluations to other languages, including English, to generalize our findings. Future research should continue to explore these ethical dimensions, ensuring that the development and evaluation of TTS systems are conducted with respect for the diversity and rights of all participants involved.

### v2

@@ 6.2 Evaluations using MUSHRA-DG
Does our proposed variant help mitigate judgement ambiguity while rating? In Table 5, we show the effects of presenting detailed guidelines along with scoresheets to 20 participants to systematically arrive at MUSHRA scores. We find MUSHRA-DG scores to be rank-consistent with the CMOS tests , while systems scores are much higher and closer to the “Excellent” label, as expected. More importantly, the standard deviation of scores across all systems reduced by 43% in Hindi and 53% in Tamil when compared to the original MUSHRA, indicating that our proposed variant is able to reduce the ambiguity of rating naturalness on a single bar while preserving ranks.
It is important to explicitly address the apparent discrepancy in the ranking of VITS and FS2 systems in Tamil to clearly validate our claim that our proposed test also preserves rankings. While the MUSHRA-DG scores for these systems in Tamil suggest a reversed ranking between these systems, a closer inspection at the scores reveals that their perceived naturalness is highly comparable. This observation is also corroborated by the CMOS scores presented in Table 3, which show minimal separation between the two systems in Tamil. To confirm this more appropriately, we conducted a focused CMOS test that directly compared FS2 and VITS. Using a scale from +3 (indicating FS2 is significantly better) to -3 (indicating VITS is significantly better), 15 listeners rated the systems, and we obtained a CMOS score of 0.13. This result reinforces the notion that the two systems perform similarly.
While system rankings in closely matched scenarios can be debated, we argue that such cases highlight the need to focus on fine-grained differences. Understanding the specific contexts in which one model outperforms another provides valuable insights into system behavior and guides targeted improvements. Keeping this in mind, we subsequently discuss the prowess of the MUSHRA-DG test in fault isolation.
@@
Similarly, for Tamil, as detailed in Figure 9 in Appendix A.5, the marginal differences across dimensions clarify the perceived inconsistency in the rankings of VITS and FS2 mentioned earlier, which, upon closer inspection, is not truly an inconsistency but rather a reflection of their comparable overall performance. We also take this opportunity to address the inflated anchor scores observed in the MUSHRA-DG test compared to the original MUSHRA scores in Tamil. One explanation for this inflation could be attributed to leniency in the scoring formula applied to subjective dimensions, resulting in higher anchor scores. However, a more compelling explanation is that the original MUSHRA test may be influenced by a Range Equalizing Bias Zielinski (2016), where participants tend to stretch scores across the entire scale. Consequently, low-quality anchors are often relegated to the extreme lower end of the scale, even when the perceived degradation is not as severe. In contrast, fine-grained evaluations like MUSHRA-DG emphasize detailed attributes and systematic scoring, which mitigate this bias. This approach reduces the tendency to artificially stretch scores, leading to higher and arguably more accurate anchor ratings. Crucially, this inflation does not affect the relative rankings of systems or the conclusions drawn from MUSHRA-DG. By minimizing variance and enhancing scoring consistency, MUSHRA-DG continues to demonstrate its robustness as a reliable and effective diagnostic and ranking tool for TTS evaluation.
Moreover, while the Reference scores appear higher in MUSHRA-DG, this increase is not merely a case of inflation. Notably, the scores of the systems also increase, resulting in all systems being classified within the “Excellent” category. This trend aligns with the expectations set by the CMOS scores shown in Table 3. Furthermore, MUSHRA-DG brings the scores of the top-performing models closer to the Reference scores, a pattern consistent with the results from CMOS evaluations. This alignment highlights the effectiveness of MUSHRA-DG in providing fine-grained assessments that capture subtle distinctions between systems while ensuring the overall scores accurately represent system performance.
Time Complexity of MUSHRA-DG. We hypothesize that the additional detail of evaluating each audio sample across multiple dimensions inevitably increases the time required for participants to complete the test. To verify this hypothesis, we visualize the average time taken across pages in Figure 6 and find that the MUSHRA-DG test indeed takes nearly twice as much time as the original MUSHRA test. However, we believe this extra time results in a much more comprehensive understanding of TTS system performance, and allows for fine-grained fault isolation, making the trade-off worthwhile.
@@ 7 Conclusion
Our comprehensive study reveals significant shortcomings in the current use of the MUSHRA test for evaluating modern high-quality TTS systems. Through an extensive analysis involving 246,000 human ratings, we identified two primary issues: reference-matching bias and judgment ambiguity. To address these issues, we propose two refined variants of the MUSHRA test: MUSHRA-NMR, which omits explicit identification of the human reference, and MUSHRA-DG, which uses detailed guidelines to calculate MUSHRA scores systematically. Our findings indicate that both variants lead to more reliable evaluations, with MUSHRA-DG offering the additional benefit of fine-grained fault isolation during assessment. Through this work, we also release MANGO, a large human rating dataset, to further support research in this area.
@@ A.2 Sensitivity of MUSHRA in Tamil
Figure 8: Spearman rank correlation of MUSHRA scores in Tamil
A.3 Visualizing MUSHRA Distributions
A.4 Detailed Guidelines for MUSHRA-DG
@@
$S_{M}=\frac{L+VQ+R}{3}-\min(MP,15)\times 5-\min(SP,7)\times 10-US\times 5-DA\times 5$
$\quad\quad\quad-WS\times 25-SEF\times 5$
We believe this formulation ensures a systematic approach to scoring, accounting for both perceptual qualities and penalizing observable errors effectively.
A.5 Fault Isolation in Tamil
Similar to Figure 5, we visualize the granular ratings across the objective and perceptual dimensions for Tamil in Figure 9.
(a) Objective (Tamil)
(b) Perceptual (Tamil)
A.6 Limitations
Our study focuses on human evaluations for Hindi and Tamil, representing a major Indo-Aryan and Dravidian language, respectively. However, we did not extend our analysis to English, a widely spoken and diverse language. This limitation is due to the scope of our current research and resource constraints. We also wish to state that while the demographics table summarizes the age and gender of participants, details for 21 out of the 492 participants are missing due to system errors; this constitutes a minor fraction and we believe does not significantly affect the analysis. While this work focuses on identifying two key issues — reference-matching bias and judgment ambiguity — several other biases may also exist in TTS evaluations, which warrant further investigation. Future studies should include evaluations in English to generalize our findings across different language families and understand how language-specific characteristics might influence TTS evaluation outcomes. This broader analysis could provide more comprehensive insights into the applicability and robustness of our proposed MUSHRA variants across diverse linguistic contexts.
A.7 Ethical Considerations
We prioritized ethical conduct throughout our research. The 492 human listeners involved in the study provided informed consent before participating in the evaluation, recruited through professional data annotation agencies. These agencies verified participant language proficiency for task relevance. We established an education criterion of completing grade 12 (Indian system) to ensure participants’ ability to accurately annotate audio content. Participants were compensated fairly for their time and expertise, following industry standards. They were also fully informed about the study nature, procedures, and their right to withdraw at any point without consequence. We ensured that our study adhered to ethical guidelines by obtaining approval from the Institutional Ethics Committee, which reviewed our methodology and confirmed compliance with ethical standards.
We release the evaluations dataset, which includes 246,000 human ratings, under CC-BY-4.0 license after careful consideration of privacy and ethical use. Identifiable information about the participants was anonymized to protect their privacy. We encourage the use of this dataset for advancing TTS evaluation metrics, emphasizing that it should be used responsibly and ethically, adhering to principles of transparency and fairness. Finally, we acknowledge that our study focuses on Hindi and Tamil, and we recognize the importance of extending such evaluations to other languages, including English, to generalize our findings. Future research should continue to explore these ethical dimensions, ensuring that the development and evaluation of TTS systems are conducted with respect for the diversity and rights of all participants involved.
@@ A.7 Ethical Considerations
Figure 12: Visualization of the MUSHRA score distributions per utterance across three systems— FS2, ST2, and VITS, along with the reference (REF) and anchor (ANC) for Tamil.